# Sparse Gaussian Neural Processes

### v1

@@ 6 Discussion
We have presented a new type of model that can be equivalently interpreted as a neural process that is amenable to GP-prior elicitation, or a meta-learner that rapidly implements sparse variational GP inference at test-time. Our model is much more computationally efficient than existing GP-based meta-learners, yet much more data efficient and interpretable than existing neural processes. This means it is particularly well suited to problems in which some combination of the following properties hold: 1. the number of datapoints in each task is beyond the GP-feasible limit, 2. domain knowledge is available in GP-prior form, 3. model interpretability is desirable, and 4. the number of observed tasks is small. Our synthetic-data experiments highlighted such utility with the fact that the (s-Conv)SGNP performance remains strong when a very small meta-dataset ($\Xi=5$) is used, while the ConvGNP performance degrades significantly. We attribute the poorer performance of the ConvSGNPs relative to the (s-Conv)SGNPs to known difficulties with Hensman et al.-style SVGP loss-landscapes (Bauer et al., 2016). In the Tétouan city power consumption experiment where all four of the above properties hold, the SGNP is by far the most performant meta-learner and the only one which allows us to imbue our prior knowledge. Furthermore, the SGNP outperforms the SPGR in the extrapolation scenario—we suspect this is because meta-learning kernel parameters across tasks allows it to learn hyperparameters than generalise better.
@@
Appendix A Experimental Details
A.1 Synthetic 1D Regression
@@ A.3 Data Pooling
A.4 Tétouan City Power Consumption
@@
Appendix B Further Discussion
Validity as a Neural Process.
A neural process encodes a context set into a representation, and then decodes the combination of the representation and the test points into a probabilistic prediction. In our model, the inducing inputs $\mathbf{Z}$ and the inducing output parameters $\mathbf{m}$ and $\mathbf{S}$ serve as the context set representation. Using the Bayesian machinery for conditioning Gaussian processes, the test points are combined with the representation to obtain posterior predictive distributions. The inference networks then serve as the neural process encoder, and we have a nonparametric decoder.
Computational Complexity.
[…]
tab:nsf_comp DeepSet ConvDeepSet Transformer $\mathcal{O}(n)$ $\mathcal{O}(n+t)$ $\mathcal{O}(n^{2})$ GP T-SVGP SVGP $\mathcal{O}(n^{3})$ $\mathcal{O}(nm^{2})$ $\mathcal{O}(bm^{2}+m^{3})$ NP ANP/TNP PT-TNP $\mathcal{O}(n_{c}+n_{t})$ $\mathcal{O}(n_{c}^{2}+n_{c}n_{t})$ $\mathcal{O}(mn_{c}+mn_{t})$ ConvGNP ConvSGNP SGNP $\mathcal{O}(n_{c}+t)$ $\mathcal{O}(t+n_{c}m^{2}{\color[rgb]{1,0,0}+n_{c}^{2}})$ $\mathcal{O}(n_{c}m^{2}{\color[rgb]{1,0,0}+n_{c}^{2}})$

### v2

@@ 6 Discussion
We have presented a new type of model that can be equivalently interpreted as a neural process that is amenable to GP-prior elicitation, or a meta-learner that rapidly implements sparse variational GP inference at test-time. Our model is much more computationally efficient than existing GP-based meta-learners, yet much more data efficient and interpretable than existing neural processes. This means it is particularly well suited to problems in which some combination of the following properties hold: 1. the number of datapoints in each task is beyond the GP-feasible limit, 2. domain knowledge is available in GP-prior form, 3. model interpretability is desirable, and 4. the number of observed tasks is small. Our synthetic-data experiments highlighted such utility with the fact that the (s-Conv)SGNP performance remains strong when a very small meta-dataset ($\Xi=5$) is used, while the ConvGNP performance degrades significantly. We attribute the poorer performance of the ConvSGNPs relative to the (s-Conv)SGNPs to known difficulties with Hensman et al.-style SVGP loss-landscapes (Bauer et al., 2016). In the Tétouan city power consumption experiment where all four of the above properties hold, the SGNP is by far the most performant meta-learner and the only one which allows us to imbue our prior knowledge. Furthermore, the SGNP outperforms the SPGR in the extrapolation scenario—we suspect this is because meta-learning kernel parameters across tasks allows it to learn hyperparameters than generalise better. Further discussion can be found in Appendix A.
@@
Appendix A Further Discussion
Validity as a Neural Process.
Our model can be viewed as a member of the conditional NP family. A conditional NP encodes a context set into a deterministic representation, and then decodes the combination of the representation and the test points into a probabilistic prediction. In our model, the inducing inputs $\mathbf{Z}$ and the inducing output parameters $\mathbf{m}$ and $\mathbf{S}$ serve as the context set representation. Using the Bayesian machinery for conditioning Gaussian processes, the test points are combined with the representation to obtain posterior predictive distributions. The inference networks then serve as the neural process encoder, and we have a nonparametric decoder. Passing the latent variable posterior parameters directly to the decoder instead of a latent variable sample is also what Volpp et al. (2021) do to convert a latent variable NP to a conditional one.
Alternatively, our model can be interpreted as a neural process with both a conditional and latent path, somewhat akin to in the original ANP (Kim et al., 2019). The inducing inputs $\mathbf{Z}$ serve as the representation in the conditional path since they are deterministic in nature, while the inducing outputs $\mathbf{u}$ serve as the probabilistic latent variable. The key innovation of our approach is to place a GP prior over the latent variable (rather than the typical spherical Gaussian). Inference over the latent variable is then performed either in closed form via the Titsias equations or in amortised fashion via a ConvDeepSet. An important difference between the SGNP’s latent path and a traditional NP’s latent path is that the predictive distribution is obtained by analytic marginalisation of the latent variable. In the non-conjugate case of the ConvSGNP, Monte Carlo sampling of the latent variable is required as normal to approximate the predictive density.
Computational Complexity.
[…]
tab:nsf_comp DeepSet ConvDeepSet Transformer $\mathcal{O}(n)$ $\mathcal{O}(n+t)$ $\mathcal{O}(n^{2})$ GP SGPR SVGP $\mathcal{O}(n^{3})$ $\mathcal{O}(nm^{2})$ $\mathcal{O}(bm^{2}+m^{3})$ NP ANP/TNP PT-TNP $\mathcal{O}(n_{c}+n_{t})$ $\mathcal{O}(n_{c}^{2}+n_{c}n_{t})$ $\mathcal{O}(mn_{c}+mn_{t})$ ConvGNP ConvSGNP SGNP $\mathcal{O}(n_{c}+t)$ $\mathcal{O}(t+n_{c}m^{2}{\color[rgb]{1,0,0}+n_{c}^{2}})$ $\mathcal{O}(n_{c}m^{2}{\color[rgb]{1,0,0}+n_{c}^{2}})$
Model Properties.
In LABEL:tab:properties we highlight which properties are exhibited by our model (SGNP) in comparison with GPs, sparse GPs (SGPR and SVGP), and the rest of the NP family. Further explanations of what we mean by each property decsription are as follows:
  1. 1.
    Prior elicitation: can we manually elicit prior beliefs (i.e. without training on synthetic/prior data)?
  2. 2.
    Task-efficient: does the method avoid the need for copious training tasks?
  3. 3.
    Scalable: is the method applicable for tasks with many observations?
  4. 4.
    Fast adaptation: does the method avoid the need for optimisation loops to make predictions on new tasks?
  5. 5.
    Interpretable: are the method’s predictions interpretable/explainable?
  6. 6.
    Modelling flexibility: can the method recover arbitrarily complex stochastic process posteriors, such as non-Gaussian posteriors or those with non-closed-form covariance structure? For example, consider the classic NP task of pixelwise regression over facial images—NPs can learn an implicit “face kernel”, but GPs are restricted to whatever closed-form kernel we choose (e.g. SE, weakly periodic etc).
This table highlights the fact that our method essentially swaps the ability to model arbitrarily complex stochastic processes with the ability to manually select the stochastic process family. Perhaps the main takeaway of our paper is that total flexibility leads to overfitting when we have limited tasks, whereas the ability to guide an NP via prior elicitation leads to greater task-efficiency.
[…]
tab:properties Property GP Sparse GP NP family SGNP Prior elicitation Task-efficient Scalable Fast adaptation Interpretable Modelling flexibility
Table 5: Properties of the various types of model considered in this work.
Appendix B Experimental Details
The code for this project can be found at https://github.com/Sheev13/sgnp.
B.1 Synthetic 1D Regression
@@ B.3 Data Pooling
B.4 Tétouan City Power Consumption